OpenStudy (anonymous):

Need help on a derivative question Compute d2y/dx2 at the point (3,3) x^2+y^2=18

OpenStudy (abb0t):

\(\sf \color{red}{2x+2y \frac{dy}{dx}=0 \Rightarrow \frac{dy}{dx} = - \frac{x}{y}}\) now take the derivative again. Can you do that?

OpenStudy (anonymous):

Ya so then the derivative of that is \[\frac{ x }{ y^2\frac{ dy }{ dx } }\]

OpenStudy (anonymous):

Correct?

OpenStudy (abb0t):

Not quite. You can use either product rule or quotient rule.

OpenStudy (anonymous):

Oh thats what I missed so then \[(-1)(y^-1)+(-x)(-y^-2)\]

OpenStudy (anonymous):

Then \[\frac{ -1 }{ y }+\frac{ x }{ y^2 }\]

OpenStudy (anonymous):

And then \[\frac{ 1+x }{ y^2 }\]

OpenStudy (anonymous):

Correct?